OpenStudy (anonymous):

f(x)=1/(x+2) f^-1(x)= (1/x)-2 Find (f^-1of)(x)

OpenStudy (anonymous):

Note that \((f^{-1}\circ f)(x) = f^{-1}(f(x))\). Since \( f^{-1}(x) = \dfrac{1}{x}-2 \), we see that \(f^{-1}(f(x)) = \dfrac{1}{f(x)}-2\). Now substitute \(f(x)=\dfrac{1}{x+2}\) into this and simplify. What do you get when you do this? :-)

OpenStudy (anonymous):

f−1(f(x)) = 1/(1/(x+2))−2. I am not too sure how to simplify this further

OpenStudy (anonymous):

Hint: 1/(1/(x+2)) = x+2. So what do you get now?

OpenStudy (anonymous):

oh, x+2-2=x so the 1's just cancel each other out?

OpenStudy (anonymous):

POP, POP!

OpenStudy (anonymous):

@Dylan548 Yup, that's correct. Does that make sense?

OpenStudy (anonymous):

Yes it does, I very much appreciate your help!!